OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 1, 3, 9, 27, ...

OpenStudy (anonymous):

Time 3?

OpenStudy (anonymous):

times*

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

my options are an = 1 • 3n + 1 an = 3 • 1n - 1 an = 1 • 3n - 1 an = 1 • 3n

OpenStudy (anonymous):

Multiply by 3 every time is what it looks like so the last one. That's what I see

OpenStudy (anonymous):

Oh ok. Thanks!